bewellplus.gsu.edu
GloriaGunzenhause650
GloriaGunzenhause650 GloriaGunzenhause650
  • 02-10-2015
  • Biology
Answered

If you flipped a penny 99 times and each time it came up heads, then the chance of the penny coming up heads again is ____ . 50% 1% 100% 25%

Verified Answer :

Kalahira
Kalahira Kalahira
  • 06-09-2017

The chance of coming up heads 99 times in a row is abysmally small: 50%^99 = 1.57 * 10^-30.
This however has already happened, it does not influence the next penny flip.
A penny has two sides, so the chance of it landing on one side is 1/2 = 0.5 = 50%.
So the result of the next flip has a chance of 50% of coming up heads.

Answer Link

Related Questions

The lighthouse is 76feet tall and has 99 stairs An ocean liner is 3 times as tall as the lighthouse What is the extra information if you want to find the height
what is name of first man made element
The American counterculture that emerged in the mid1960s was marked by
what is the expanded form for 7026
How do the rods and cones in the eye differ A Rods are very sensitive to light but do not distinguish colors BCones are less sensitive to light but can distingu
An atom has no electrical charge because
Identifying Adjectives and the Words They Modify Identify the adjectives and the words they modify in the following sentences Do not include a an or the EXAMPLE
How did big bend national park get its name
Read the summary of Thomas Paines Common Sense 1 In Common Sense Thomas Paine attempts to appeal to the colonists 2 He argues that America must separate from Br
What is characteristic of George Grenville